# The Brownian loop measure on Riemann surfaces and applications to length spectra

### v1

@@ 1 Introduction
The Brownian loop measure in the complex plane $\mathbb{C}$ was introduced in [25, 22] for the study of Schramm–Loewner evolutions and has various applications to random conformal geometry, see, e.g., [43, 27, 1, 12, 19, 11]. The definition can be extended to any Riemannian surface $(X,g)$ with or without boundary, and we denote it by $\mu_{X}$:
$\mu_{X}:=\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\int_{X}\mathbb{W}_{x\to x}^{t}(X)\,\mathrm{d}\mathrm{vol}_{g}(x),$ (1.1)
where $\mathbb{W}_{x\to x}^{t}(X)$ is a Brownian bridge on $X$ from $x\in X$ to $x$ of total quadratic variation $4t$ with respect to the metric $g$, killed upon hitting the boundary of $X$, and whose total mass is the heat kernel $p_{X}(x,x;t)$ with Dirichlet boundary condition. See Section 2 for more details.
The Brownian loop measure is an infinite but sigma-finite measure on unrooted and oriented loops in $X$ that satisfies two important properties: conformal invariance, and the restriction property as shown in [25, 1], thanks to the weight $1/t$ in the integral (1.1). We recall the proof in Theorem 2.2. Therefore, we consider $\mu_{X}$ as a measure on oriented loops on a Riemann surface instead of a Riemannian surface (as only the conformal class of the background metric $g$ matters). Throughout the paper, all loops are oriented unless otherwise specified. To understand how Brownian loops are distributed on a Riemann surface, we first compute the Brownian loop measure in each free homotopy class of closed curves.
Let $(X,g)$ be a Riemannian surface without boundary (or we replace $X$ with boundary by $X\setminus\partial X$). From the uniformization theorem, if $X$ is not conformally equivalent to one of ${}^\hat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, and tori, then $X$ admits a unique complete hyperbolic metric that is conformal to $g$ and may have cusps and funnels. We say a free homotopy class of oriented closed curves on $X$ is not homotopic to a point, if it is not a trivial homotopy class nor a class homotopic to a cusp. Every such homotopy class of oriented closed curves on $X$ contains a unique hyperbolic geodesic $\gamma$. We define the iteration number $m(\gamma)\in\mathbb{Z}_{>0}$ if $\gamma$ is $m(\gamma)$-time iteration of a primitive closed geodesic.
@@
If $(X,g)$ admits a complete hyperbolic metric that is conformal to $g$, the total mass of the Brownian loop measure in a homotopy class of closed curves on $(X,g)$ that is not homotopic to a point equals
@@
It follows immediately that the total mass of Brownian loops in a homotopy class of a simple closed geodesic on $X$ is invariant if we change the conformal structure of $X$ by twisting along the same geodesic. Our proof follows from a standard computation using the heat kernel on the hyperbolic plane $\mathbb{H}$. The same result for closed surfaces (i.e., compact and without boundary) was stated briefly in [28, Sec. 7] using [30].
When $X$ is one of $\hat{\mathbb{C}}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, all homotopy classes are homotopic to a point. The only remaining case is when $X$ is a torus.
@@
We note that although the Brownian loop measure is conformally invariant and a priori does not prefer any particular metric in the same conformal class, Lemmas 1.1 and 1.2 show that it does single out the length of the geodesic for the hyperbolic metric (or the Euclidean metric if $X$ is a torus).
@@ 1.1 Identity on surface length spectra
Let $X$ be a complete hyperbolic surface without boundary (perhaps with cusps or funnels). Let $\{p_{1},\cdots,p_{n},\cdots\}\subset X$ be a closed subset consisting of finitely or countably many distinct points and $X^{\prime}=X\setminus\{p_{1},\cdots,p_{n},\cdots\}$ be the Riemann surface with the complex structure induced from $X$. Then $X^{\prime}$ admits a unique complete hyperbolic metric conformal to the complex structure. On $X^{\prime}$, the points $p_{n}$ are cusps. We obtain the following identity between the length spectra of $X$ and $X^{\prime}$.
@@
We note that this identity holds regardless of the punctures’ number or location, and there are infinitely many (but countable) terms on the right-hand side. A similar result holds when $X$ is a torus. See Theorem 4.3. The proof of Theorem 1.3 uses Lemma 1.1 and the fact that the Brownian loop measure is conformally invariant and that a countable family of points is polar (i.e., almost surely never hit by a Brownian motion).
@@
$\sum_{\gamma}\frac{1}{\mathrm{e}^{\ell_{\gamma}(X)}+1}=\frac{1}{2}$
where the summation is taken over all (unoriented) simple closed geodesics. It would be interesting to consider whether Theorem 1.3 and the McShane identity have any relationship.
To our best knowledge, not so many types of identity on the geodesic lengths on hyperbolic surfaces are known. McShane identity [31] and Mirzakhani’s generalization [32] are established by decomposing horocycle (or boundary) according to the homotopy class of geodesic flow starting from the boundary and ending when intersecting the boundary or self-intersecting. Basmajian identity [2] is established by decomposing the boundary according to the homotopy class of geodesic flow with both two endpoints on the boundary. See also [36] for a unification and generalization of McShane identity and Basmajian identity. Bridgeman identity [5, 6, 10, 9] is established by decomposing the unit tangent bundle according to the homotopy class of geodesic flow with both two endpoints on the boundary. Luo–Tan identity [29] is established by decomposing the unit tangent bundle according to the homotopy class of geodesic flow ending when self-intersect in both two directions. See [7] for a survey and the references therein about these four types of identity and their generalizations.
@@ 1.2 Explicit computations of Brownian loop measure
We note that Lemma 1.1 covers the case where $(X,g)$ has a non-empty boundary. After removing the boundary, $X$ has a unique complete hyperbolic metric conformal to $g$. One may think of the boundary as being “infinitely far”, and their components are funnels. This allows us to compute the Brownian loop measure using the geodesic length in many setups. We select a few simple examples here.
@@
Here, $\gamma^{m}$ is the curve obtained by iterating $m$ times the circle $e^{-r/2}S^{1}$, which is the unique simple closed geodesic on $\mathbb{A}_{r}$ endowed with the complete hyperbolic metric and has length $2\pi^{2}/r$, in the counterclockwise direction. See Figure 1.
Lawler also computed this quantity in the proof of [23, Prop. 3.9] using the Brownian bubble measure. See Remark 4.5. The total mass of Brownian loops of odd winding number was also used in [24].
Another immediate consequence is the following relation between the electrical thickness [20] and Brownian loop measure. Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 00 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 00 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 00 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$, where $\mathbb{D}=\{z\in\hat{\mathbb{C}}\,|\,|z|<1\}$ and $\mathbb{D}^{*}=\{z\in\hat{\mathbb{C}}\,|\,|z|>1\}$. The electrical thickness of $K$ is defined as
@@
This function is nonnegative, thanks to the Grunsky inequality. See, e.g., [39, Thm. 4.1, (21)]. Restricting $\theta(\cdot)$ to the universal Teichmüller curve $\mathcal{T}(1)$ (which is a natural fiber space over the universal Teichmüller space $T(1)$), $\theta(\cdot)$ was shown [47, Ch. I, Thm. 5.3] to be a Kähler potential for the unique homogeneous Kähler metric (Velling–Kirillov metric) on $\mathcal{T}(1)$.
@@
$\mu_{\hat{\mathbb{C}}}(\{\text{loops with winding $m$ around }0\text{ intersecting }K\})=\frac{\theta(K)}{2\pi^{2}m^{2}}+\frac{1}{2|m|}.$
We mention that the Kähler potential (i.e., the Loewner energy [48] / universal Liouville action [47]) on the connected component $T_{0}(1)$ of $T(1)$ for the unique homogeneous Kähler metric (Weil–Petersson metric) also has an expression in terms of the Brownian loop measure [49]. Investigating the relation between these two Kähler potentials in terms of the Brownian loop measure will be interesting.
@@
Assume that $X$ is a Riemann surface of finite type (i.e., the Euler characteristic is finite) and that $X$ cannot be obtained by removing finitely many (including zero) points from a closed Riemann surface. Then, the total mass of Brownian loops on $X$ that are not homotopic to a point is finite.
It is not hard to construct an infinite type surface $X$ that has at least one funnel and infinitely many geodesics of length less than $1$ (e.g., by gluing infinitely many pairs of pants). This implies that $X$ has an infinite mass of noncontractible Brownian loops and explains the assumption on the finite type of $X$. We also note that the total mass of Brownian loops not homotopic to a point is infinite if $X$ is a closed Riemann surface, hence also if $X$ is obtained by removing finitely many (we already assumed finite type) points from a closed surface (by Corollary 4.1). Now, we discuss how to renormalize the total mass of the Brownian loop measure on a closed hyperbolic surface.
@@ 1.3 Zeta-regularized determinant of Laplacian
We note that the integral $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}N_{X}(L)$ is formally the total mass of Brownian loops homotopic to a primitive geodesic by Lemma 1.1, although infinite. This result expresses $\log\mathrm{det}_{\zeta}\Delta$ of a closed hyperbolic surface $X$ only in terms of its genus (since Area$(X)=2\pi(2\operatorname{genus}(X)-2)$ by the Gauss–Bonnet theorem) and a rather simple function of the length spectrum of $X$. Moreover, the normalizing integral with respect to $\tilde{\mathop{\rm Li}}(\mathrm{e}^{L})$ is universal (does not depend on $X$).
@@ 2 Brownian loop measure on Riemannian surfaces
In this section, we recall the definition and basic properties of the Brownian loop measure on a Riemannian surface $(X,g)$.
Let $x\in X$, $t>0$, we denote by $\mathbb{W}^{t}_{x}(X)$ the sub-probability measure on the path of the Brownian motion run at speed $2$ on $X$ (i.e., the diffusion of the infinitesimal generator $-\Delta_{g}$ where $\Delta_{g}=-\operatorname{div}_{g}\operatorname{grad}_{g}$ is the positive Laplace–Beltrami operator) starting from $x$ on the time interval $[0,t]$, which is killed upon hitting $\partial X$ (therefore $\mathbb{W}^{t}_{x}(X)$ has a total mass less than $1$ if $X$ has a non-polar boundary). The measures $\mathbb{W}^{t}_{x\to y}$ on paths from $x$ to $y$ are obtained from disintegrating $\mathbb{W}^{t}_{x}(X)$ according to the endpoint $y$:
$\mathbb{W}^{t}_{x}(X)=\int_{X}\mathbb{W}^{t}_{x\to y}(X)\,\mathrm{d}\mathrm{vol}_{g}(y).$
@@
Note that $\mathbb{W}^{t}_{x\to y}(X)$ has total mass $p_{X}(x,y;t)$, where $p_{X}$ is the heat kernel on $(X,g)$ (and with Dirichlet boundary condition if $\partial X\neq\emptyset$). By our convention, the total quadratic variation of a path under the measure $\mathbb{W}_{x\to y}^{t}$ is almost surely $4t$.
Define the Brownian loop measure on $X$:
$\mu_{X}:=\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\int_{X}\mathbb{W}_{x\to x}^{t}(X)\,\mathrm{d}\mathrm{vol}_{g}(x).$
Since we integrate the family of measures on paths with the same starting and ending points, it is an infinite measure on the set of loops. We will forget the starting point and the time parametrization; however, we will keep the loop’s orientation. So we consider $\mu_{X}$ as a measure on unparametrized but oriented loops.
@@
  • (Restriction property) If $X^{\prime}\subset X$, then $\mathrm{d}\mu_{X^{\prime}}(\eta)=\mathbf{1}_{\eta\subset X^{\prime}}\,\mathrm{d}\mu_{X}(\eta)$, where $\mathbf{1}_{\cdot}$ is the indicator function.
@@
modulo time parametrization. See [1, Lem. 3.1] or [25, Sect. 3.2.1].
Now, we define the rooted loop measure

### v2

@@ Abstract
Keywords: Brownian loop measure, Riemann surface, geodesic length, identity of length spectrum, determinant of Laplacian
MSC 2020: 30F60, 57M50, 60J65
@@ 1 Introduction
The Brownian loop measure in the complex plane $\mathbb{C}$ was introduced in [27, 24] for the study of Schramm–Loewner evolutions and has various applications to random conformal geometry, see, e.g., [46, 29, 1, 12, 21, 11]. The definition can be extended to any orientable Riemannian surface $(X,g)$ with or without boundary, and we denote it by $\mu_{X}$. We recall its definition in Section 2.
The Brownian loop measure $\mu_{X}$ is an infinite but sigma-finite measure on the space of loops in $X$, built from the Brownian motion, that satisfies two important properties: conformal invariance, and the restriction property as shown in [27, 1]. We recall the proof of this fact in Theorem 2.2. Therefore, $\mu_{X}$ is considered as a measure on the space of loops on a Riemann surface instead of a Riemannian surface (as only the conformal class of the background metric $g$ matters). Throughout the paper, all loops are oriented. To understand how Brownian loops are distributed on a Riemann surface, we first compute the Brownian loop measure in each free homotopy class of closed curves.
Let $(X,g)$ be a Riemannian surface without boundary (or we replace $X$ with boundary by $X\setminus\partial X$). From the uniformization theorem, if $X$ is not one of ${}^\widehat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, and tori, then $X$ admits a unique complete hyperbolic metric that is conformal to $g$. In particular, $X$ may have cusps and funnels and may be of infinite type. We say a free homotopy class of oriented closed curves on $X$ is essential if it is not a trivial homotopy class nor a class homotopic to a cusp. Every such homotopy class of oriented closed curves on $X$ contains a unique hyperbolic geodesic $\gamma$. We define the iteration number $m(\gamma)\in\mathbb{Z}_{>0}$ if $\gamma$ is $m(\gamma)$-time iteration of a primitive closed geodesic.
@@
If $(X,g)$ admits a complete hyperbolic metric that is conformal to $g$, the total mass of the Brownian loop measure in an essential homotopy class of closed curves on $(X,g)$ equals
@@
It follows immediately that the total mass of Brownian loops in a homotopy class of a simple closed geodesic on $X$ is invariant if we change the conformal structure of $X$ by twisting along the same geodesic. Our proof follows from a standard computation using the heat kernel on the hyperbolic plane $\mathbb{H}$. The same result for closed surfaces (i.e., compact and without boundary) was obtained in [31] (see also [30, Sec. 7]) using [33].
When $X$ is one of $\widehat{\mathbb{C}}$, $\mathbb{C}$, $\mathbb{C}\setminus\{0\}$, all homotopy classes are not essential. The only remaining case is when $X$ is a torus.
@@
In this case, the flat metrics on $T$ are not unique, but are related by a scaling which leaves ${\rm Area}(T)/(\pi\ell_{\gamma}(T)^{2})$ invariant.
We note that although the Brownian loop measure is conformally invariant and a priori does not prefer any particular metric in the same conformal class, Lemmas 1.1 and 1.2 show that it does single out the length of the geodesic for the complete hyperbolic metric (or the flat metrics if $X$ is a torus).
@@ 1.1 Identity on surface length spectra
Let $X$ be a complete hyperbolic surface without boundary (perhaps with cusps or funnels). Let $P=\{p_{1},\cdots,p_{n},\cdots\}\subset X$ be a closed subset consisting of finitely or countably many distinct points and $X^{\prime}=X\setminus\{p_{1},\cdots,p_{n},\cdots\}$ be the Riemann surface with the complex structure induced from $X$. Then $X^{\prime}$ admits a unique complete hyperbolic metric compatible with the complex structure. On $X^{\prime}$, the points $p_{n}$ are cusps. We obtain the following identity between the length spectra of $X$ and $X^{\prime}$.
@@
We note that this identity holds regardless of the punctures’ number or location, and there are infinitely many (but countable) terms on the right-hand side. A similar result holds when $X$ is a torus. See Theorem 4.3.
Remark 1.4.
The proof of Theorem 1.3 uses Lemma 1.1 and the fact that the Brownian loop measure is conformally invariant and that a countable family of points is polar (i.e., almost surely never hit by a Brownian motion). In fact, Theorem 1.3 also holds when we replace $P$ by any closed, polar, subset of $X$.
@@
$\sum_{\gamma}\frac{1}{\mathrm{e}^{\ell_{\gamma}(X)}+1}=1$
where the summation is taken over all oriented simple closed geodesics. It would be interesting to consider whether Theorem 1.3 and the McShane identity have any relationship.
To our best knowledge, not so many types of identity on the geodesic lengths on hyperbolic surfaces are known. McShane identity [34] and Mirzakhani’s generalization [35] are established by decomposing horocycle (or boundary) according to the homotopy class of geodesic flow starting from the boundary and ending when intersecting the boundary or self-intersecting. Basmajian identity [2] is established by decomposing the boundary according to the homotopy class of geodesic flow with both two endpoints on the boundary. See also [39] for a unification and generalization of McShane’s identity and Basmajian identity. Bridgeman identity [5, 6, 10, 9] is established by decomposing the unit tangent bundle according to the homotopy class of geodesic flow with both two endpoints on the boundary. Luo–Tan identity [32] is established by decomposing the unit tangent bundle according to the homotopy class of geodesic flow ending when self-intersect in both two directions. See [7] for a survey and the references therein about these four types of identity and their generalizations.
@@ 1.2 Explicit computations of Brownian loop measure
We note that Lemma 1.1 covers the case where $(X,g)$ has a non-empty boundary. After removing the boundary, $X$ has a unique complete hyperbolic metric conformal to $g$. One may think of the boundary as being “infinitely far”. This allows us to compute the Brownian loop measure using the geodesic length in many setups. We select a few simple examples here.
@@
Here, $\gamma^{m}$ is the curve obtained by iterating $m$ times the circle $e^{-r/2}\mathbb{S}^{1}$ counterclockwise, which is the unique simple closed geodesic on $\mathbb{A}_{r}$ endowed with the complete hyperbolic metric and has length $2\pi^{2}/r$, in the counterclockwise direction. See Figure 1.
Lawler also computed this quantity in the proof of [25, Prop. 3.9] using the Brownian bubble measure. See Remark 4.5. The total mass of Brownian loops of odd winding number was also used in [26].
Another immediate consequence is the following relation between the electrical thickness [22] and Brownian loop measure. Let $K\subset\mathbb{C}\setminus\{0\}$ be a connected compact set separating 00 and $\infty$. Let $\Omega$ and $\Omega^{*}$ be respectively the connected component of $\mathbb{C}\setminus K$ containing 00 and $\infty$. Let $f$ be a conformal map $\mathbb{D}\to\Omega$ fixing 00 and $h$ a conformal map $\mathbb{D}^{*}\to\Omega^{*}$ fixing $\infty$, where $\mathbb{D}=\{z\in\widehat{\mathbb{C}}\,|\,|z|<1\}$ and $\mathbb{D}^{*}=\{z\in\widehat{\mathbb{C}}\,|\,|z|>1\}$. The electrical thickness of $K$ is defined as
@@
This function is nonnegative, thanks to the Grunsky inequality. See, e.g., [42, Thm. 4.1, (21)]. Restricting $\theta(\cdot)$ to the universal Teichmüller curve $\mathcal{T}(1)$ (which is a natural fiber space over the universal Teichmüller space $T(1)$), $\theta(\cdot)$ was shown [50, Ch. I, Thm. 5.3] to be a Kähler potential for the unique homogeneous Kähler metric (Velling–Kirillov metric) on $\mathcal{T}(1)$.
@@
$\mu_{\widehat{\mathbb{C}}}(\{\text{loops with winding $m$ around }0\text{ intersecting }K\})=\frac{\theta(K)}{2\pi^{2}m^{2}}+\frac{1}{2|m|}.$
We mention that the Kähler potential (i.e., the Loewner energy [51] / universal Liouville action [50]) on the connected component $T_{0}(1)$ of $T(1)$ for the unique homogeneous Kähler metric (Weil–Petersson metric) also has an expression in terms of the Brownian loop measure [52]. It would be interesting to investigate the relation between these two Kähler potentials in terms of the Brownian loop measure.
@@
Assume that $X$ is a Riemann surface of finite type (i.e., the Euler characteristic is finite) and that $X$ has infinite hyperbolic area. Then, the total mass of Brownian loops on $X$ that are in essential homotopy classes is finite.
The condition on $X$ is equivalent to being a Riemann surface of finite type that cannot be obtained by removing finitely many (including zero) points from a closed Riemann surface and therefore has at least one funnel.
It is not hard to construct an infinite type surface $X$ that has at least one funnel and infinitely many geodesics of length less than $1$ (e.g., by gluing infinitely many pairs of pants with boundary lengths less than $1$). This implies that $X$ has an infinite mass of noncontractible Brownian loops and explains the assumption on the finite type of $X$. We also note that the total mass of all Brownian loops in essential homotopy classes is infinite if $X$ is a closed Riemann surface (see Theorem 4.10), hence also infinite if $X$ is obtained by removing finitely many points from a closed surface (by Corollary 4.1). Now, we discuss how to renormalize the total mass of the Brownian loop measure on a closed hyperbolic surface.
@@ 1.3 Zeta-regularized determinant of Laplacian
We note that the integral $\int_{L=0}^{\infty}\frac{1}{\mathrm{e}^{L}-1}\ \mathrm{d}N_{X}(L)$ is formally the total mass of Brownian loops homotopic to a primitive geodesic by Lemma 1.1, although infinite. This result expresses $\log\mathrm{det}_{\zeta}\Delta$ of a closed hyperbolic surface $X$ only in terms of its genus (since Area$(X)=2\pi(2\operatorname{genus}(X)-2)$ by the Gauss–Bonnet theorem) and a rather simple function of the length spectrum of $X$. Moreover, the normalizing integral with respect to $\widetilde{\mathop{\rm Li}}(\mathrm{e}^{L})$ is universal (does not depend on $X$).
@@ 2 Brownian loop measure on Riemannian surfaces
In this section, we recall the definition and basic properties of the Brownian loop measure on a Riemannian surface $(X,g)$. For readers not familiar with Brownian motion on Riemannian manifolds, we have included some background in Appendix A.
Let $x\in X$, $t>0$, we denote by $\mathbb{W}^{t}_{x}(X)$ the sub-probability measure on the path of the Brownian motion, in the space $C^{0}([0,t],X)$ endowed with the Borel $\sigma$-algebra induced from the uniform convergence topology, run at speed $2$ on $X$ starting from $x$ on the time interval $[0,t]$, which is killed upon hitting $\partial X$. In other words, $\mathbb{W}^{t}_{x}(X)$ has total mass
$|\mathbb{W}_{x}^{t}(X)|=\mathbb{W}_{x}(X)(\{W\,|\,\text{which is well defined on }[0,t]\})\leq 1,$
where $\mathbb{W}_{x}(X)$ is the probability distribution of the Brownian motion on $X$ starting from $x$, whose diffusion generator is the Laplace–Beltrami operator $\Delta=\operatorname{div}_{g}\operatorname{grad}_{g}$. By our convention, the total quadratic variation of a path under the measure $\mathbb{W}_{x}^{t}$ is almost surely $4t$.
We now define the measures $\mathbb{W}^{t}_{x\to y}(X)$ on $C^{0}([0,t],X)$ on paths from $x$ to $y$ by disintegrating $\mathbb{W}^{t}_{x}(X)$ according to the endpoint $y\in X$:
$\mathbb{W}^{t}_{x\to y}(X):=\lim_{\varepsilon\to 0+}\frac{\mathbb{W}^{t}_{x}(X)\mathbf{1}_{\{W_{t}\in D_{\varepsilon}(y)\}}}{\mathrm{vol}_{g}(D_{\varepsilon}(y))}$
where $D_{\varepsilon}(y)$ is the metric ball of radius $\varepsilon$ around $y\in X$ and $\mathbf{1}_{\cdot}$ is the indicator function.
@@
In particular, we have the equality of measures
$\mathbb{W}^{t}_{x}(X)=\int_{X}\mathbb{W}^{t}_{x\to y}(X)\,\mathrm{d}\mathrm{vol}_{g}(y),$
and $\mathbb{W}^{t}_{x\to y}(X)$ has total mass
$\left\lvert\mathbb{W}^{t}_{x\to y}(X)\right\rvert:=\lim_{\varepsilon\to 0+}\frac{\mathbb{W}^{t}_{x}(X)(\{W\in C^{0}([0,t],X)\,|\,W_{t}\in D_{\varepsilon}(y)\})}{\mathrm{vol}_{g}(D_{\varepsilon}(y))}=p_{X,g}(x,y;t),$
where $p_{X,g}$ is the heat kernel on $(X,g)$ from the definition of the Brownian motion with Dirichlet boundary condition (see Appendix A).
The parametrized Brownian loop measure on $X$ is defined as the measure on $\{W\in C([0,t],X)\,|\,t\geq 0,W_{0}=W_{t}\}$,
$\mu_{X}^{*}:=\int_{0}^{\infty}\frac{\mathrm{d}t}{t}\int_{X}\mathbb{W}_{x\to x}^{t}(X)\,\mathrm{d}\mathrm{vol}_{g}(x).$
Since we integrate the family of measures on paths with the same starting and ending points, it is an infinite measure on the set of parametrized loops.
The Brownian loop measure is the induced measure $\mu_{X}$ on the equivalence classes where $W\in C^{0}([0,t],X)\sim\widetilde{W}\in C^{0}([0,s],X)$ if and only if there is an increasing continuous time-reparametrization of loops $i:[0,t]_{/0\sim t}\to[0,s]_{/0\sim s}$ such that $\widetilde{W}_{i(r)}=W_{r}$ for all $r$. Namely, $\mu_{X}$ is the measure obtained from $\mu_{X}^{*}$ by forgetting the starting point and the time parametrization, but remembering the orientation. The topology induced on the equivalent classes of loops is the uniform sup-norm $d_{X}^{\operatorname{loop}}$ on distance in $X$ modulo increasing reparametrization, and $\mu_{X}$ is Borel for this topology.
@@
  • (Restriction property) If $X^{\prime}\subset X$, then $\mathrm{d}\mu_{X^{\prime}}(\eta)=\mathbf{1}_{\eta\subset X^{\prime}}\,\mathrm{d}\mu_{X}(\eta)$.
@@
modulo time parametrization. See [1, Lem. 3.1] or [27, Sect. 3.2.1].
Now, the rooted loop measure is written as